# Imitation Learning for Combinatorial Optimisation under Uncertainty

### v1

@@ 33    $i\leftarrow i+1$;
34      Train policy $\pi_{i+1}$ on dataset $\mathcal{D}$;
35
@@ 34      Train policy $\pi_{i+1}$ on dataset $\mathcal{D}$;
Algorithm 1 General DAgger Algorithm
1Function VanillaDAgger($x_{k},\hat{a}_{ik},\pi_{i}$):
2 $a_{k}\leftarrow\pi_{i}(x_{k})$;
3 $\beta_{i}\leftarrow\lambda^{i-1}\beta_{0}$;
4 $z\sim\text{Uniform}(0,1)$;
5
6 if $z\leq\beta_{i}$ then
7    return $\hat{a}_{ik}$;
8
9 else
10    return $a_{k}$;
11
12   end if
13
14 end
Algorithm 2 VANILLA DAgger Decision Rule
Expert Type.
The set of experts $\mathcal{E}$ may include different model types, such as Myopic, Deterministic, Full-Information, and Two-Stage Stochastic experts. Multi-stage stochastic experts are typically excluded due to their high computational cost.
Expert Optimality.
The optimality of expert actions depends on the complexity of the underlying SDP. As such problems often involve large combinatorial decision spaces, approximation techniques are commonly employed. For example, a Two-Stage Stochastic expert may be restricted to a limited number of scenarios to maintain tractability. The parameter $\kappa$ specifies the expert stopping criterion and may represent a time limit, a target optimality gap, or other termination conditions or even a combination of multiple criteria.
Expert Interaction.
Experts may be queried either once or iteratively. In a one-shot setting, expert interaction is limited to a single iteration ($I=1$), after which the policy is trained offline. In interactive settings, the stopping criterion allows the algorithm to perform $I$ iterations, resulting in a sequence of learned policies $\{\pi_{1},\dots,\pi_{I}\}$.
Labelling Strategy.
The function $F_{j}$ determines the number of expert calls $J$ made at a given state. It may also skip expert queries for selected states, for instance when querying is expensive or when certain states provide limited learning value. Selective querying strategies have been explored in the literature [zhang2016query]. If an expert returns multiple actions in a single call, then ${a}^{*}_{ijk}$ represents a vector of actions. The aggregation function $A$ combines these actions into a single target label.
Number of Experts.
When $|\mathcal{E}|>1$, the algorithm operates with multiple experts. The function $F_{e}$ determines which expert is queried for each of the $J$ calls at a given state.
Trajectory Control.
The decision rule (DR) specifies which action is executed at each decision epoch. Many ML-based approaches rely on full expert control, executing expert actions throughout the trajectory. A common alternative is the Vanilla DAgger rule shown in Algorithm˜2, which executes the expert action with probability $\beta_{i}$ and the learner’s action with probability $1-\beta_{i}$. The probability $\beta_{i}$ typically decreases over iterations as a function of $\lambda$ and $\beta_{0}$, allowing the learner to gradually transition from expert-guided behaviour to autonomous decision making.
7 Results
@@
To demonstrate the practical effectiveness of Algorithm˜1, we evaluate its performance on the physician-to-patient assignment (PPA) problem. All expert optimisation models are solved using Gurobi. The full experimental codebase is publicly available at GitHub. All experiments are conducted on a machine equipped with an Intel Xeon Gold 6258R CPU running at 2.70 GHz. The remainder of this section is organised as follows. Section˜7.1 describes the experimental setup, including instance generation, evaluation metrics, baseline policies, and the computational environment. Section˜7.2 presents and discusses the experimental results in detail, while Section˜7.3 draws inferences from these results.
7.1 Experimental Setup
@@
We begin by describing the parameters used to generate instances of the PPA. The proposed approach is evaluated on synthetically generated instances constructed under controlled conditions. The total number of patient calls arriving at the clinic, denoted by $K$, is sampled from a normal distribution with mean $\mu=100$ and standard deviation $\sigma=8$. Each patient is assigned a priority class: class 1 (high priority) with probability $0.3$, and class 2 (regular priority) otherwise.
To capture the tendency of high-priority patients to call later in the day, we generate an arrival score for each patient from a beta distribution, $\mathrm{Beta}(a,b)$. Specifically, arrival scores are drawn with parameters $(a=3,b=1)$ for class 1 patients and $(a=1,b=1)$ for class 2 patients. Patients are then ordered according to these arrival scores, ensuring that high-priority patients are more likely to appear towards the end of the scheduling horizon.
@@
The second baseline, denoted by $\pi^{t}$, corresponds to a two-stage stochastic expert. At each decision epoch, this policy samples 30 future demand scenarios and invokes the expert solver to determine an action.
@@
The final baseline is the exact optimal policy, denoted by $\pi^{*o}$, computed separately for each test instance. This policy represents an upper bound on achievable performance and serves as a benchmark for assessing the quality of learned and heuristic policies.
7.2 Performance Evaluation
@@
We evaluate multiple configurations of expert models using Algorithm˜1 and compare their performance against the baseline policies. All methods are tested on a dataset of 1,000 instances. For each policy, we report the average objective value, the total number of rejected patients by priority class, and the total number of assignments to non-preferred physicians, aggregated across both priority classes. Each expert model is given a time limit of 30 seconds to compute an action at each decision epoch.
7.2.1 Baselines
@@
Table˜3 summarises the performance of the baseline policies. The greedy policy $\pi_{\text{greedy}}$ performs poorly, primarily due to a high number of rejections among priority 1 patients. In contrast, the two-stage stochastic policy $\pi^{t}$ achieves substantially better performance by reducing rejections across both priority classes. This improvement comes at the cost of a modest increase in assignments to non-preferred physicians.
@@
It is important to note that $\pi^{t}$ is computationally expensive. With 30 sampled future scenarios, the policy requires approximately 25 hours to complete 10 episodes when executed in parallel. $\pi^{*o}$ shows the best achievable performance with significantly low priority 1 patient rejections.
Policy Cost Priority 1 Rejected Priority 2 Rejected Undesirable Assignments
$\pi_{greedy}$ 3047.48 10.26 15.26 22.39
$\pi^{t}$ 1785.58 2.86 19.98 27.42
$\pi^{*o}$ 1271.25 0.45 20.45 25.32
Table 3: Average performance metrics for $\pi_{greedy}$, $\pi^{t}$ and $\pi^{*o}$
7.2.2 Learning without interaction (Isolated Learning)
@@
We first report results obtained without expert interaction. The hyperparameter settings used for Algorithm˜1 are summarised in Table˜4 and correspond to the base configuration of the algorithm. Any deviations from this configuration are explicitly stated in the relevant experimental subsections. Setting $I=1$ implies that only a single outer iteration is performed, resulting in the training of a single learned policy.
We consider three expert policies: the full-information expert $\pi^{*o}$, the deterministic expert $\pi^{*p}$, and the two-stage stochastic expert $\pi^{*t}$. Each expert is queried once per state. To construct the dataset of expert demonstrations, we simulate a total of $H=4000$ episodes. Due to the imposed 30-second time limit on expert optimisation, only the two-stage stochastic expert $\pi^{*t}$ may terminate early $(\textbf{{MIP}}_{\textbf{SPPA}})$ and return a non-optimal action. As a result, the learned policy is ultimately trained using demonstrations generated by $\pi^{t}$. In contrast, both the full-information and deterministic experts can solve their respective optimisation models ($\textbf{{MIP}}_{\textbf{PPA}}$) to optimality within the allotted time.
The use of a time limit is necessary to keep the overall training time tractable. Since only a single learning iteration is performed, expert actions are selected using the vanilla DAgger decision rule. We also note that employing the full-information expert $\pi^{*o}$ within Algorithm˜1 is equivalent to solving deterministic instances of the PPA using a mixed-integer programming solver. Finally, to accelerate data collection, 10 episodes are simulated in parallel.
@@
Table 4: Hyper Parameters for Algorithm˜1
Full-Information and Deterministic Experts
We set $F_{e}(\cdot)$ to $\pi^{*o}$ and $\pi^{*p}$ for the full-information and deterministic expert configurations, respectively. Using Algorithm˜1, we collect approximately 400,000 expert demonstrations. Solving the $\textbf{{MIP}}_{\textbf{PPA}}$ requires, on average, 0.2 seconds per instance.
We first analyse the effect of the number of simulated episodes on the performance of the learned policy, to understand how the size of the demonstration dataset influences learning quality. The results are shown in Figure˜4. For the full-information expert, increasing the number of demonstrations does not lead to a clear or consistent improvement in performance. In contrast, policies trained using demonstrations from the deterministic expert exhibit a steady performance improvement as the number of demonstrations increases, ultimately achieving superior results.
A plausible explanation is that the deterministic expert encounters a wider range of system states, as it does not select actions that are optimal with respect to the realised trajectory. Consequently, the learned model is exposed to a more diverse state distribution, including states that would not typically be visited under an optimal policy. This increased diversity improves the generalisation capability of the learned policy when deployed. The black line in Figure˜4 represents the performance of the baseline policy $\pi^{t}$.
@@
Table˜5 reports detailed results for the best-performing policies learned using the optimal full-information and deterministic experts. The higher cost relative to $\pi^{t}$ is primarily attributable to an increased number of rejected priority 1 patients under the policy learned from the full-information expert. The table also reports the total training time required to simulate 4,000 episodes.
@@
It is important to note that the full-information expert can label all states within an episode in a single optimisation run by solving the $\textbf{{MIP}}_{\textbf{PPA}}$ once, and therefore does not require the use of Algorithm˜1 for data generation. In contrast, the deterministic expert must generate and solve a new optimisation instance at every state within an episode, which substantially increases the overall data collection time.
Learnt Policy Cost

### v2

@@ 29    $i\leftarrow i+1$;
30    $\hat{\pi}_{i+1}\leftarrow\textsc{Train}(\mathcal{D})$;
31
@@ 30    $\hat{\pi}_{i+1}\leftarrow\textsc{Train}(\mathcal{D})$;
Algorithm 3 Adaptive DAgger Algorithm
We characterize the imitation learning framework with respect to the taxonomy: the expert type is given by the set $\mathcal{E}$, which may include myopic, deterministic, full-information, or two-stage stochastic models; expert optimality, governed by a stopping criterion $\kappa$ that controls solution accuracy and computational effort; and expert interaction, determined by the number of iterations $I$, yielding a sequence of policies $\{\hat{\pi}_{1},\dots,\hat{\pi}_{I}\}$. Additional design choices include the labelling strategy $F_{j}$, which specifies when and how often expert queries are issued and how multiple actions $a^{*}_{ijk}$ are aggregated via $A$; the number of experts $|\mathcal{E}|$, with expert selection handled by $F_{e}$; and trajectory control, defined by a decision rule $DR$.
6 Application to Generalised Assignment Problem
This section presents the problem setting and experimental framework used to evaluate the proposed imitation learning methodology. We first introduce the dynamic physician-to-patient assignment problem, which serves as a representative sequential decision problem under uncertainty and allows for systematic instantiation of the expert taxonomy. We then describe the experimental design, including expert configurations, learning setups, and baseline policies, and report a comprehensive empirical evaluation of the proposed framework. The experiments are designed to assess how different expert types, levels of optimality, and interaction strategies influence learning performance, data efficiency, and computational cost.
6.1 Generalised Assignment Problem
The classical Generalised Assignment Problem (GAP) has been extensively studied since the seminal work of [balachandran1976integer]. The problem consists of assigning a set of items (or jobs) to a set of agents (or machines) such that each job is assigned to exactly one agent, agent capacity constraints are respected, and the total assignment cost or profit is optimised. GAP has found applications in a wide range of domains, including job scheduling [balachandran1976integer], routing [fischetti2019machine], and patient admission scheduling [ceschia2011local]. A standard integer programming formulation of the GAP is given by (8)-(11). Here, $c_{ij}\in\mathbb{R}_{+}$ denotes the cost incurred by assigning job $j$ to agent $i$, $d_{ij}\in\mathbb{R}_{+}$ is the resource consumption of agent $i$ when processing job $j$, $L_{i}\in\mathbb{R}_{+}$ is the capacity of agent $i$, and $x_{ij}$ is a binary decision variable equal to $1$ if job $j$ is assigned to agent $i$ and 0 otherwise.
$\min\quad$ $\sum_{i=1}^{m}\sum_{j=1}^{n}c_{ij}x_{ij}$ (8)
s.t. $\sum_{i=1}^{m}x_{ij}=1,\quad$ $\forall j=1,\dots,n,$ (9)
$\sum_{j=1}^{n}d_{ij}x_{ij}\leq L_{i},\quad$ $\forall i=1,\dots,m,$ (10)
$x_{ij}\in\{0,1\},\quad$ $\forall i=1,\dots,m,\ \forall j=1,\dots,n.$ (11)
6.1.1 Physician-to-Patient Assignment
Appointment requests arrive independently over time and are handled individually by a receptionist. At each arrival, the receptionist must decide whether to accept or reject the request based on the current state of the system. If the request is accepted, the receptionist must assign the patient to a physician. Each patient $k$ has a preferred physician $p_{k}$ among a set of eligible physicians $\mathcal{P}_{k}\subseteq\mathcal{P}$. Each patient is characterised by a service duration $d_{k}\in\mathbb{R}_{+}$, which is estimated at the time of the call, and a priority class $r_{k}\in\{1,2\}$. The arrival process is such that calls from priority 1 patients arrive near the end of the horizon.
Rejecting a patient of priority $r_{k}$ incurs a penalty $c^{\text{rej}}_{r_{k}}$, with $c^{\text{rej}}_{1}>c^{\text{rej}}_{2}$. In addition, assigning a patient to an undesired but eligible physician incurs a preference penalty $c^{\text{pref}}_{r_{k}}$. The primary source of uncertainty in the system arises from the random arrival times and sequences of patients within a session. This creates a trade-off between accepting lower-priority patients early and preserving capacity for potential high-priority patients who may arrive later. Furthermore, assigning patients to undesired physicians may reduce future capacity, potentially forcing rejections later in the booking horizon. The resulting problem is an SDP in which each decision epoch corresponds to a patient arrival, and the action is to assign the patient to an eligible physician or to reject the request. The offline mixed-integer programming formulation, denoted by $(\textbf{{MIP}}_{\textbf{PPA}})$, is provided in the appendix.
We adopt the PPA problem to demonstrate the expert taxonomy. It requires anticipating future arrivals and strategically reserving capacity for high-priority patients, a core challenge in sequential decision making under uncertainty. This structure allows us to distinguish between myopic, deterministic, and stochastic experts based on how they incorporate future information. At the same time, the underlying CO formulations remain relatively tractable, ensuring that the computational bottleneck of repeated expert queries in DAgger is manageable. Feasibility is also straightforward to enforce through capacity constraints, enabling clear evaluation of learned policies without extensive post-processing. Finally, PPA requires rapid decision-making at each patient call-in, making it well-suited for assessing the benefits of learned policies that replace expensive optimisation with fast deployment-time decisions. We emphasise that our goal is not to propose a new solution method for the PPA itself, but rather to use it as a representative SDP for systematically analysing expert design choices within IL.
6.1.2 MDP Formulation
We now formulate the PPA problem as an MDP.
State Variable – At decision epoch $k$, the system state is given by $x_{k}=\big(x^{k}_{\text{patient}},x^{k}_{\text{physician}},x^{k}_{\text{eligibility}}\big)$. The component $x^{k}_{\text{patient}}=[d_{k},r_{k}]$ represents the characteristics of the current patient. The component $x^{k}_{\text{physician}}=[x^{k}_{p}]_{p\in\mathcal{P}}$ captures the status of all physicians, where $x^{k}_{p}=\big({1}_{p=p_{k}},x^{k}_{pl},x^{k}_{p1},x^{k}_{pw}\big).$ Here, ${1}_{p=p_{k}}$ indicates whether physician $p$ is the patient preferred physician, $x^{k}_{l}$ denotes the remaining assignment capacity, $x^{k}_{p1}$ is the number of priority-1 patients already assigned, and $x^{k}_{pw}$ is the current workload. A list of indicators gives the eligibility vector functions $x^{k}_{\text{eligibility}}=[{1}_{p\in\mathcal{P}_{k}}]_{p\in\mathcal{P}}.$
Decision Variable – At each decision epoch $k$, the receptionist selects an action $a_{k}\in\mathcal{A}_{k}=\{0\}\cup\mathcal{P}_{k}$, where $a_{k}=0$ corresponds to rejecting the request and $a_{k}=p$ corresponds to assigning the patient to physician $p$. A physician $p$ is eligible only if $d_{k}+x^{k}_{pw}\leq L_{p}$, ensuring feasibility of all actions in $\mathcal{A}_{k}$.
Exogenous Information – The exogenous information at epoch $k$ corresponds to the characteristics of the next arriving patient and is denoted by $\xi_{k}=[d_{k+1},r_{k+1},p_{k+1},\mathcal{P}_{k+1}].$
Transition Function – The transition function $X^{M}(\cdot)$ maps the system from state $x_{k}$ to $x_{k+1}$ according to $x_{k+1}=X^{M}(x_{k},a_{k},\xi_{k}).$ Specifically, for the selected physician $p=a_{k}$,
  • $x^{k+1}_{pl}=x^{k}_{l}-{1}$,
  • $x^{k+1}_{p1}=x^{k}_{p1}+{1}_{r_{k}=1}$,
  • $x^{k+1}_{pw}=x^{k}_{pw}+d_{k}$.
$C_{k}(x_{k},a_{k})=\begin{cases}c^{\text{rej}}_{r_{k}},&\text{if }a_{k}=0,\\[6.0pt] c^{\text{pref}}_{r_{k}},&\text{if }a_{k}\neq 0\text{ and }a_{k}\neq p_{k},\\[6.0pt] 0,&\text{otherwise.}\end{cases}$
6.2 Results
@@
To demonstrate how different expert properties influence learning outcomes, we evaluate the performance of Algorithm˜3 on the physician-to-patient assignment (PPA) problem. All expert optimisation models are solved using Gurobi. The full experimental codebase is publicly available at GitHub. All experiments are conducted on a machine equipped with an Intel Xeon Gold 6258R CPU running at 2.70  GHz. The remainder of this section is organised as follows. Section˜6.2.1 describes the experimental setup, including instance generation, evaluation metrics, baseline policies, and the computational environment. Section˜6.2.2 presents and discusses the experimental results in detail.
6.2.1 Experimental Setup
@@
- We begin by describing the parameters used to generate instances of the PPA. The proposed approach is evaluated on synthetically generated instances constructed under controlled conditions. The total number of patient calls arriving at the clinic is denoted by $K$, as each patient corresponds to a decision epoch. $K$ is sampled from a normal distribution with mean $\mu=100$ and standard deviation $\sigma=8$. Each patient is assigned a priority class: class 1 (high priority) with probability $0.3$, and class 2 (regular priority) otherwise.
To capture the tendency of high-priority (emergency) patients to call later in the day, we generate an arrival score for each patient from a beta distribution, $\mathrm{Beta}(a,b)$. Specifically, arrival scores are drawn with parameters $(a=3,b=1)$ for class 1 patients and $(a=1,b=1)$ for class 2 patients. Patients are then ordered according to these arrival scores, ensuring that high-priority patients are more likely to appear towards the end of the scheduling horizon.
@@
The second baseline, denoted by $\pi^{t}$, corresponds to a two-stage stochastic expert. At each decision epoch, this policy samples 30 future demand scenarios. These scenarios are generated by simulating future arrivals conditioned on the current system state. The expert solver is then invoked to compute an action by solving the resulting two-stage stochastic optimisation problem over the generated scenarios.
@@
The final baseline is the exact optimal full-information policy, denoted by $\pi^{*f}$, computed separately for each test instance. This policy represents an upper bound on achievable performance and serves as a benchmark for assessing the quality of learned and heuristic policies.
6.2.2 Performance Evaluation
@@
We evaluate multiple configurations of expert models using Algorithm˜3 and compare their performance against the baseline policies. All methods are tested on a dataset of 1,000 instances. For each policy, we report the average objective value, the total number of rejected patients by priority class, and the total number of assignments to non-preferred physicians, aggregated across both priority classes. Each expert model is given a time limit of 30 seconds to compute an action at each decision epoch.
Baselines
@@
- Table˜4 summarises the performance of the baseline policies. The table reports the average cost alongwith the % of rejections for each priority and % of undesirable assignments by each policy. The main difference between the performance of all policies is due to priority 1 patient rejections. The greedy policy $\pi_{\text{greedy}}$ performs poorly, primarily due to a high percentage of rejections among priority 1 patients. In contrast, the two-stage stochastic policy $\pi^{t}$ achieves substantially better performance by reducing rejections across both priority 1 patients. This improvement comes at the cost of a modest increase in assignments to non-preferred physicians.
@@
It is important to note that $\pi^{t}$ is computationally expensive despite the time limit. With 30 sampled future scenarios, the policy requires approximately 25 hours to complete 10 episodes when executed in parallel. $\pi^{*f}$ shows the best achievable performance with significantly low priority 1 patient rejections.
Table 4: Average performance metrics for $\pi_{greedy}$, $\pi^{t}$ and $\pi^{*f}$
Policy Cost Priority 1 Rejected % Priority 2 Rejected % Undesirable Assignments (%)
$\pi_{greedy}$ 3047.48 39.7 20.2 22.3
$\pi^{t}$ 1785.58 10.4 26.4 27.5
$\pi^{*f}$ 1271.25 1.4 26.2 25.5
Learning without interaction (Isolated Learning)
@@
- We first report results obtained without expert interaction. The hyperparameter settings used for Algorithm˜3 are summarised in Table˜5 and correspond to the base configuration of the algorithm. Any deviations from this configuration are explicitly stated in the relevant experimental subsections. Setting $I=1$ implies that only a single outer iteration is performed, resulting in the training of a single learned policy.
Table 5: Hyper Parameters for Algorithm˜3
@@
$F_{e}(\cdot)$ $\pi^{*f}$, $\pi^{*d}$ $\pi^{t}$
We consider three expert policies as targets: the full-information expert ($\pi^{*f}$), the deterministic expert ($\pi^{*d}$), and the two-stage stochastic expert ($\pi^{*t}$). We cannot include the multistage expert $\pi^{*}$, as it is too difficult to obtain even for small instances of the problem. Each expert is queried once per state. The corresponding learned policies are denoted by $\hat{\pi}^{d}$, $\hat{\pi}^{*d}$, $\hat{\pi}^{*d}$. To construct the dataset of expert demonstrations, we simulate a total of $H=4000$ episodes. Using Algorithm˜3, we collect approximately 400,000 expert demonstrations for each of the three expert targets. Due to the imposed 30-second time limit on expert optimisation, only the two-stage stochastic expert $\pi^{*t}$ terminates early $(\textbf{{MIP}}_{\textbf{SPPA}})$ and may return a non-optimal action. As a result, the learned policy is ultimately trained using demonstrations generated by $\pi^{t}$. In contrast, both the full-information and deterministic experts can solve their respective optimisation models ($\textbf{{MIP}}_{\textbf{PPA}}$) to optimality within the allotted time. We also note that employing the full-information expert $\pi^{*f}$ within Algorithm˜3 is equivalent to solving deterministic instances of the PPA using a mixed-integer programming solver in one shot.
To employ the two-stage stochastic expert ($\pi^{t}$), it is necessary to specify the number of future scenarios generated at each decision epoch. In this study, we use a fixed number of scenarios across all epochs. Alternative designs could allow the number of scenarios to vary over time—for example, by generating more scenarios in later stages, where the two-stage problem may be easier to solve, and decisions are likely to be more critical with respect to the utilisation of remaining capacity. We evaluate three configurations using 5, 15, and 30 scenarios. Finally, data collection is parallelised using 10 threads, with 10 episodes simulated simultaneously. All training times and results reported in this study are based on this parallel execution setup.
We first analyse the effect of the number of simulated episodes on the performance of the learned policy, to understand how the size of the demonstration dataset influences learning quality. The results are shown in Figure˜4. For the full-information expert, increasing the number of demonstrations does not lead to a clear or consistent improvement in performance. In contrast, policies trained using demonstrations from the deterministic expert exhibit a steady performance improvement as the number of demonstrations increases.
A plausible explanation is that the deterministic expert encounters a wider range of system states, as it does not select actions that are optimal with respect to the realised trajectory. Consequently, the learned model is exposed to a more diverse state distribution, including states that would not typically be visited under an optimal policy. This increased diversity improves the generalisation capability of the learned policy when deployed. The black line in Figure˜4 represents the performance of the baseline policy $\pi^{t}$. However, learning from a two-stage stochastic expert with multiple scenarios substantially outperforms learning from a full-information and deterministic expert based on a single scenario.
@@
Table˜6 reports detailed results for the best-performing policies learned using all three experts. The higher costs of $\hat{\pi}^{*f}$, $\hat{\pi}^{*d}$ relative to $\pi^{t}$ baseline are primarily attributable to an increased number of rejected priority 1 patients under the policy learned from the full-information expert. The table also reports the total training time required to simulate 4,000 episodes. It is important to note that the full-information expert can label all states within an episode in a single optimisation run by solving the $\textbf{{MIP}}_{\textbf{PPA}}$ once, and therefore does not require the use of Algorithm˜3 for data generation. In contrast, the deterministic expert must generate and solve a new optimisation instance at every state within an episode, which substantially increases the overall data collection time.
The learned policies using ${\pi}^{t}$ as the target markedly reduce patient rejections. Among the configurations considered, training with 30 scenarios yields the best performance and comes closest to the baseline policy $\pi^{t}$, as expected. These improvements come at the cost of increased training time, since multiple scenarios must be generated and the corresponding multi-scenario optimisation problem $(\textbf{{MIP}}_{\textbf{SPPA}})$ must be solved. The results indicate substantial performance improvements as the size of the training dataset increases. Performance gains are expected to plateau once the number of episodes exceeds approximately 2,000.
Table 6: Average performance metrics for the best $\hat{\pi}^{\cdot}$
Learnt Policy Scenarios Cost
Priority 1
Rejected %
Priority 2
Rejected%
Undesirable
Assignments (%)
Train Time (hours)
Full-Information ($\hat{\pi}^{f}$) 1 2031.34 13.2 29.5 25.8 0.5
Deterministic ($\hat{\pi}^{d}$) 1 1975.72 11.8 29.8 25.5 3.25
Two-Stage ($\hat{\pi}^{t}$) 5 1890.35 12.1 26.4 27.5 8.65
15 1822.45 10.8 26.6 27.4 32.47
30 1801.85 10.4 26.8 27.2 65.94
Table˜7 reports detailed statistics on optimality gaps and solution times under the imposed 30-second time limit, aggregated over approximately 400,000 expert calls for the two-stage stochastic expert using $\textbf{{MIP}}_{\textbf{SPPA}}$. Most instances are solved to optimality or near optimality, as indicated by the 99% values. As the number of scenarios increases, a modest rise in the average optimality gap is observed. When gaps are non-zero, this is typically due to the solver identifying the optimal solution but requiring additional time to certify optimality. Overall, these results indicate that the stochastic formulation remains tractable and provides high-quality expert demonstrations.
Average solution times increase with the number of scenarios. As a result, deploying $\pi^{t}$ with 30 scenarios for real-time decision making would require approximately 5.09 seconds per decision on average, whereas the learned policy produces decisions within milliseconds. Solution times also vary over the planning horizon. As shown in Figure˜5, early decision epochs exhibit larger optimality gaps and longer solve times, reflecting the larger problem size at the beginning of an episode. Under a 30-second time limit, solution times at early epochs can approach 20 seconds on average, further underscoring the value of fast and reliable learned policies. On the other hand, solving the $\textbf{{MIP}}_{\textbf{PPA}}$ for the full-information and deterministic expert requires, on average, 0.2 seconds per instance.
Table 7: Statistics across expert ($\pi^{t}$) calls
5 0 0 0.66 6.41
15 0.004 0.16 2.51 30
30 0.013 0.29 5.09 30
Figure 5: Average Gap and Average solve time across decision epoch
Learning with interaction
Figure˜6 reports the average cost across learning iterations for all expert types. The horizontal axis corresponds to the learned policy $\pi_{i}$ obtained after iteration $i$. The results show a consistent decrease in average cost as additional data are incorporated through expert interaction. Performance metrics for the best learned policies are summarised in Table˜8. We also show the average objectives of the best models obtained through isolated learning, indicated by dotted lines in the figure.
The models learned through interaction achieve lower average costs compared with those obtained via isolated learning, as reported in Table˜6. Table˜8 also reports the percentage improvement relative to the isolated learning models. In addition, policies learned with interaction reject fewer patients overall. No significant performance difference is observed between policies learned from the full-information expert and those learned from the deterministic expert.
As shown in Figure˜6, learning performance begins to saturate after approximately 200 iterations, although the best-performing models are obtained at later iterations. We would like to point out that the best models obtained through isolated learning required 4000 training episodes. Using interaction, the learned model improves on those values within 2000 episodes. This highlights the effect of interaction. Compared with the full-information and deterministic expert settings, a stronger and more consistent decreasing trend in total cost is observed as learning progresses. Among all configurations, the policy trained using 30 scenarios achieves the best performance. It yields a marginal improvement over the policy learned without interaction, while requiring a significantly smaller number of expert demonstrations, coming very close to the baseline policy $\pi^{t}$.
Training times are also reported in Table˜8. Under interactive learning, the training time associated with the deterministic expert more than doubles. For the full-information expert, training time increases substantially due to the use of the vanilla DAgger decision rule. Under this rule, the states encountered during learning differ from those observed by a pure full-information expert, which necessitates solving the $\textbf{{MIP}}_{\textbf{PPA}}$ at each newly encountered state. Similarly, the two-stage stochastic expert also sees a considerable increase in training times.
Figure˜6(b) shows the same plot with respect to training time in minutes. Within the first 6 hours of training, we see that the Two-Stage Stochastic Expert with 5 scenarios shows better test performance than those with 15 and 30. Similarly, the one with 15 scenarios is better than 30 scenarios for the next 10 hours. This suggests adding more states and learning more efforts in the short term is better. But eventually, if one allows more time, the one with higher scenarios gives better models.
Figure 6: Test performance of different learned models
(a) With respect to learned models
(b) With respect to time
[…]
Table 8: Average performance metrics for the best model with interaction
Expert Scenarios Best Model Cost % Improvement
Priority 1
Rejected%
Priority 2
Rejected%
Undesirable
Assignments (%)
Train Time (hours)
Full-Information ($\hat{\pi}^{f}$) 1 261 1951.76 3.9 10.9 29.8 26.9 16.40
Deterministic ($\hat{\pi}^{d}$) 1 371 1943.85 1.6 11.3 29.3 25.6 14.77
Two-stage Stochastic ($\hat{\pi}^{t}$) 5 389 1844.75 2.4 10.65 27.4 27.2 23.19
15 301 1804.59 0.9 10.1 27.2 27.0 59.22
30 251 1795.93 0.3 10.0 27.1 27.2 101.83
Comprising Optimality of Experts
Figure˜6 shows that the two-stage stochastic expert achieves the best performance, albeit at the cost of substantial training times. To reduce the overall computational burden, we investigate whether training time can be lowered by relaxing the optimality requirement of the two-stage stochastic expert. As illustrated in Figure˜5, the first few decision epochs are the most computationally demanding and account for the majority of total training time. In comparison, a fixed 30-second time limit appears overly conservative for later epochs, where problem instances are smaller and easier to solve.
To address this imbalance, we introduce an additional stopping criterion based on the MIP gap. Specifically, we set the tolerance to $2\%$, meaning that the expert terminates once this optimality gap is reached. As shown in Figure˜5, the observed gaps are, on average, already well below this threshold. We therefore set $\kappa=2\%$ and repeat the same experiment using the two-stage stochastic expert under this relaxed optimality condition.
Table˜9 reports the best policies obtained across all three scenario settings using the two-stage stochastic expert. The learned policies match the performance of those in Table˜8, which were trained under a fixed 30-second time limit. Notably, the policy trained with 30 scenarios achieves even stronger results than its counterpart in Table˜8. One possible explanation is that solving the two-stage problem to near optimality as suggested by the MIP gaps obtained, may bias the expert targets toward the sampled scenarios. In contrast, terminating based on an MIP gap criterion may yield targets that better approximate the underlying multistage stochastic optimisation problem represented by $\pi^{*}$. We also observe a substantial reduction in training times—particularly for the 15- and 30-scenario settings—without any deterioration in policy performance.
Table 9: Average performance metrics for the best model with interaction with MIP gap 0.2%
Expert Scenarios Best Model Cost